### v1

@@ 3.2.2 Output Generation
Using equations (5.3) and (2), we generate one-month-ahead forecasts for fire count and burnt area for all time points after the initial 36-month window. Given XGBoost’s high capacity for fitting complex patterns, robust cross-validation is essential to prevent overfitting. We adopt a Super Learner-style cross-validation scheme (Van der Laan et al.,, 2007) for both training and evaluation. The training data span 2011–2022, and 2023 is held out as the test set. Since temporal information is already embedded in the window-based format, random partitioning does not compromise temporal dependencies.
[…]
@@ 3.3 Stage II: Bayesian Latent Gaussian Modelling
We incorporate the one-month-ahead forecasts of fire count and burnt area from equations (5.3) and (2) as external covariates in a Bayesian latent Gaussian model, which yields the final probabilistic predictions. This two-stage setup reflects the principles of stacked generalisation (Wolpert,, 1992), where the predictions from one model serve as informative inputs to a second, more interpretable model to enhance overall performance.
@@ 3.3.1 Model Structure and INLA
and this joint posterior could be approximated using MCMC. However, under the latent Gaussian model framework, it suffices to compute the marginal posterior distributions $\pi(\eta_{i}\mid\boldsymbol{y})$ and $\pi(\theta_{j}\mid\boldsymbol{y})$ for further inference of $\boldsymbol{y}$, thanks to the conditional independence assumption. Rue et al., (2009) proposed a fast approximation algorithm for marginal posteriors in latent Gaussian models using the integrated nested Laplace approximation (INLA). This algorithm was later reformulated to accommodate big data settings, enhancing numerical scalability and enabling fast inference (Van Niekerk et al.,, 2023). These developments form the foundation of the R-INLA package (Martins et al.,, 2013).
@@
with integration points $\boldsymbol{\theta}_{k}$ and weights $\Delta_{k}$. The marginal posterior of the linear predictors $\pi(\eta_{i}\mid\boldsymbol{y})$ is derived in a similar manner. Staring with the Gaussian approximation $\widetilde{\pi}(\boldsymbol{u}\mid\boldsymbol{\theta},\boldsymbol{y})$, the conditional posterior $\pi(\eta_{i}|\boldsymbol{\theta},\boldsymbol{y})$ is then also approximated as Gaussian. Its mean and variance can be efficiently computed leveraging the linear relationship $\boldsymbol{\eta}=\boldsymbol{Au}$ and some tricks on computing the $i$-th diagonal element of the inverse of the precision matrix associated with $\widetilde{\pi}(\boldsymbol{u}|\boldsymbol{\theta},\boldsymbol{y})$ (Van Niekerk et al.,, 2023). The marginal posterior $\pi(\eta_{i}|\boldsymbol{y})$ is then approximated by:
@@
Finally, Van Niekerk et al., (2023); van Niekerk and Rue, (2024) proposed a low-rank correction to the mean of $\widetilde{\pi}(\boldsymbol{u}\mid\boldsymbol{\theta},\boldsymbol{y})$ using variational Bayes, further improving the approximation of both $\widetilde{\pi}(u_{i}\mid\boldsymbol{y})$ and $\widetilde{\pi}(\eta_{i}\mid\boldsymbol{y})$.
@@ 3.3.2 Likelihood Specification
Although Poisson and Tweedie likelihoods perform well in modelling general wildfire trends via XGBoost, they are not optimal in the latent Gaussian framework. Specifically, the Tweedie distribution, being composed of i.i.d. Gamma components, features a light right tail and is therefore unsuitable for capturing extreme burnt areas. Instead, we adopt the extended generalised Pareto (eGP) distribution (Naveau et al.,, 2016) for modelling the burnt area. This distribution combines features of Gamma and Pareto distributions by applying a power transformation to the standard Pareto distribution. It is defined on the positive real line and behaves like a Gamma for small values and like a Pareto for large values. The right tail behaviour is controlled by a shape parameter $\xi$, allowing flexibility in modelling heavy tails.
@@
We define three linear predictors for each council $s$ at time $t$: $\eta^{Z}_{s,t}$, $\eta^{C}_{s,t}$ and $\eta^{B}_{s,t}$, corresponding to fire presence ($Z$), fire count ($C$) and burnt area ($B$), respectively. The full hierarchical model is described as:
$Z_{s,t}\mid\eta_{s,t}^{Z}$ $\sim\text{Bernoulli}\{\text{logit}^{-1}(\eta_{s,t}^{Z})\}$
$\{Y_{s,t}^{C}\mid\eta_{s,t}^{C},Z_{s,t}=1\}$ $\sim\text{Trucated Poisson}\{\exp(\eta_{s,t}^{C})\}$
$\left\{\sqrt{Y_{s,t}^{B}}\mid\eta_{s,t}^{B},Z_{s,t}=1\right\}$ $\sim\text{eGP}\{\sigma\left(\eta_{s,t}^{B}\right),\xi,\kappa\}$
@@ 3.3.3 Effects in the Linear Predictor
Although the XGBoost predictions $\widehat{y}^{C}_{s,t}$ and $\widehat{y}^{B}_{s,t}$ could be included as fixed effects, we avoid this approach as it imposes a linear relationship with the linear predictor. Given the complexity introduced by spatial and spatio-temporal interactions, such an assumption is overly restrictive. Instead, we model the effect of $\widehat{y}^{(\cdot)}_{s,t}$ assuming a first-order random walk (RW1) prior on 20 discretised bins of the prediction $\widehat{y}_{s,t}^{(\cdot)}$:
$R_{k}-R_{k-1}\sim N(0,\tau^{-1}_{R}),$ (6)
where $R_{k}$ represents the effect of the $k$-th bin of the discretised covariate and $\tau_{R}$ is the precision parameter. As the fire count and burnt area contribute to the Poisson and eGP likelihoods only when fire is present (i.e., when $y^{C}_{s,t}>0$) in the hurdle model training, we discretise $\widehat{y}_{s,t}^{C}$ and $\widehat{y}_{s,t}^{B}$ conditional on ${y}_{s,t}^{C}>0$ when constructing the $R_{k}$ in their respective linear predictors $\eta_{s,t}^{C}$ and $\eta_{s,t}^{B}$. By contrast, $\widehat{y}_{s,t}^{C}$ and $\widehat{y}_{s,t}^{B}$ are discretised unconditionally when constructing the corresponding $R_{k}$ in $\eta_{s,t}^{Z}$.
We incorporate spatially structured and unstructured effects through the Besag–York–Mollié model, using the reparameterised BYM2 formulation (Simpson et al.,, 2017), as implemented in R-INLA. A BYM2 effect $\boldsymbol{b}$ combines a scaled intrinsic conditional autoregressive (CAR) component $\boldsymbol{\delta}$ (with unit variance) and unstructured noise $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$, where $\boldsymbol{I}$ represents the identity matrix with dimension defined by the length of $\boldsymbol{\epsilon}$, as
@@
$\eta_{s,t}^{Z}$ $=\beta_{0}^{Z}+$ $G_{c}(s,t;\tau_{G_{c}},\phi_{G_{c}})+$ $G_{d}(s,t;\tau_{G_{d}},\phi_{G_{d}})+T(t;\tau_{T}^{Z})+R(\widehat{y}_{s,t}^{C};\tau_{R}^{Z^{C}})+R(\widehat{y}_{s,t}^{B};\tau_{R}^{Z^{B}}),$ (7)
$\eta_{s,t}^{C}$ $=\beta_{0}^{C}+\beta_{1}^{C}$ $G_{c}(s,t;\tau_{G_{c}},\phi_{G_{c}})+\beta_{2}^{C}$ $G_{d}(s,t;\tau_{G_{d}},\phi_{G_{d}})+T(t;\tau_{T}^{C})+R(\widehat{y}_{s,t}^{C};\tau_{R}^{C}),$
$\eta_{s,t}^{B}$ $=\beta_{0}^{B}+\beta_{1}^{B}$ $G_{c}(s,t;\tau_{G_{c}},\phi_{G_{c}})+\beta_{2}^{B}$ $G_{d}(s,t;\tau_{G_{d}},\phi_{G_{d}})+T(t;\tau_{T}^{B})+R(\widehat{y}_{s,t}^{B};\tau_{R}^{B}).$
@@ 3.3.4 Priors
We now elaborate on the priors used for the eGP likelihood hyperparameters ($\kappa$ and $\xi$) as well as those associated with the linear predictor. For $\kappa$ and $\xi$, we adopt Penalised Complexity (PC) priors following the framework of Simpson et al., (2017), which provide a principled mechanism to control model complexity by penalising deviations from a simpler base model. This deviation is measured via the Kullback–Leibler divergence (KLD; Kullback and Leibler, 1951), and the corresponding distance is defined as $d=\sqrt{2\text{KLD}}$. An exponential prior is then placed on this distance, yielding a memoryless penalty on increasing model complexity. The PC prior for each parameter is obtained by transforming this exponential prior back to the original parameter scale.
In the special case of the standard GDP ($\kappa=1$), Opitz et al., (2018) derived a PC prior for $\xi>0$, using the base model with $\xi=0$. They computed the KLD between a GPD density $f_{\xi}(y)=f_{\text{GPD}}(y;\xi)$ and the base model $f_{\xi_{0}}(y)=f_{\text{GPD}}(y;\xi=0)$ as:
@@
where $\lambda$ is the rate parameter of the exponential distribution, controlling the strength of penalisation. The two priors are nearly indistinguishable for large $\lambda$ (e.g., $\lambda>3$), but diverge notably for smaller values, as illustrated in Figure 3 of Opitz et al., (2018).
@@
with derivation provided in the Supplementary Materials A.2. Following the approach for deriving a PC prior for the parameter $\xi$, we may either use the exact KLD in (10), or apply a second-order Taylor expansion around $\kappa=1$, which gives
@@
Figure 5 displays $\pi_{1}(\kappa)$ and $\pi_{2}(\kappa)$ under various values of $\lambda$. When $\lambda$ is large (e.g. $\lambda>5$), both priors concentrates around $\kappa=1$, exhibiting similar behaviour. As $\lambda$ decreases, the mode of $\pi_{1}(\kappa)$ shifts leftwards towards zero and diverge as $\kappa\rightarrow 0$, whereas $\pi_{2}(\kappa)$ remains locally symmetric around $\kappa=1$. This suggests that $\pi_{1}(\kappa)$ always shrinks $\kappa$ towards a value in $(0,1]$, and it is less suitable for expressing weakly informative priors over $\kappa>1$ compared to $\pi_{2}(\kappa)$. To determine the more appropriate prior between $\pi_{1}(\kappa)$ and $\pi_{2}(\kappa)$, we consider the interpretation and functional role of $\kappa$. According to Naveau et al., (2016), $\kappa$ governs the lower-tail behaviour of the cumulative distribution function of an eGP random variable $Y$ via
@@
Figure 5: PC priors for $\kappa$ based on exact KLD in (10) (left) and the approximated KLD around $\kappa=1$ (11) (right) under 6 penalisation rates $\lambda\in\{0.1,0.5,1,2,5,10\}$.
@@ 4.1 Model Comparison
To evaluate the effectiveness of incorporating XGBoost-based wildfire forecasts into our modelling framework and to assess the appropriateness of the eGP likelihood for wildfire modelling, we compare several variations of the latent Gaussian model differing in their likelihood choices and linear predictor components. Let M1 denote the full two-stage model introduced in Section 3. To examine the role of the eGP likelihood, we substitute it with two alternative distributions commonly used in environmental applications: the Gamma and Weibull likelihoods, yielding models M2 and M3, respectively. Additionally, we construct M4, identical to M1 but excluding the XGBoost-derived effects $R$ in $\eta^{Z}$, $\eta^{C}$, and $\eta^{B}$ (see (7)), to represent forecasting without access to future covariates, relying solely on Gaussian random and fixed effects.
We use wildfire data from 2011 to 2022 as the training set and data from 2023 as the test set. One-month-ahead forecasts of fire count and burnt area are generated as described in Section 3.2.2. The posterior predictive distributions of the fire presence $Z_{s,t}$, fire count $Y_{s,t}^{C}$ and square-root-transformed burnt area $\sqrt{Y_{s,t}^{B}}$ are obtained from 1000 posterior simulations. To evaluate performance, we use Area Under Curve (AUC) to assess the predictive accuracy for fire presence, where the posterior mean of $\widehat{Z}_{s,t}$ serves as the estimate of $\mathbb{P}(Z_{s,t}=1)$. The positive burnt area $\sqrt{\widehat{Y}_{s,t}^{B}}\mid Z_{s,t}=1$ is evaluated by continuous ranked probability score (CRPS), which is a proper scoring rule that measures the difference between a predictive distribution $F$ and a single observation $y$ by
$\text{CRPS}(F,y)=\int_{\mathbb{R}}\left[F(t)-\mathbbm{1}(t\geq y)\right]^{2}\mathrm{d}t,$
@@
Since forecasts are typically communicated as probabilities over categorised bins, with particular emphasis on large events, we also compute weighted binned scores following the structure proposed by Opitz, (2023). Specifically, for a fire count threshold vector $\boldsymbol{h}^{C}=(0,1,2,\ldots,10,15,20,25,30)$, the weighted scoring for fire count $r^{C}$ is the weighted sum of squared residuals between predicted and empirical probabilities across all observations and thresholds:
$r^{C}=\sum_{s,t}\sum_{h\in\boldsymbol{h}^{C}}w^{C}(h)\left[\widehat{\mathbb{P}}(Y_{s,t}^{C}\leq h)-\mathbbm{1}(Y_{s,t}^{C}\leq h)\right]^{2},$
@@
which increases approximately linearly with $h$. Here, $\mathbb{P}(Y_{s,t}^{C}\leq h)$ is the unconditional probability of $Y_{s,t}^{C}$ obtained by integrating out $Z_{s,t}$ from the conditional distribution $Y_{s,t}^{C}\mid Z_{s,t}=1$.
@@
$r^{B}=\sum_{s,t}\sum_{h\in\boldsymbol{h}^{B}}w^{B}(h)\left[\widehat{\mathbb{P}}(Y_{s,t}^{B}\leq h)-\mathbbm{1}(Y_{s,t}^{B}\leq h)\right]^{2},$
@@
Metric M1 (eGP) M2 (Gamma) M3 (Weibull) M4 (eGP, no XGB)
AUC 0.883 0.881 0.881 0.791
CRPS 4.70 4.77 4.66 4.92
Unweighted $r^{C}$ 350 349 347 371
Weighted $r^{C}$ 5.05 4.74 4.65 4.91
Unweighted $r^{B}$ 549 558 550 586
Weighted $r^{B}$ 14.5 14.6 14.0 14.5
Table 1 summarises the model performance on the test set in 2023 across six evaluation metrics. Among models M1, M2, and M3, which differ only in the likelihood, no single likelihood dominates across all metrics. Model M1, using the eGP likelihood, achieves the highest AUC (0.883) and the lowest unweighted burnt area score (549). It is better than M2 (Gamma likelihood) for burnt area modelling, since the right tail of eGP has been adapted to allow for a heavy tail. However, its fire count performance lags slightly behind M2. Model M3 (Weibull likelihood), which can also accommodate heavy-tailed behaviour depending on its shape parameter, achieves the best CRPS, unweighted and weighted fire count score, and weighted burnt area score. M3 attains the best performance in four out of six metrics, whereas M1 leads in two. Nonetheless, performance differences between M1, M2, and M3 are marginal, and reasons for this will be further discussed in Section 5.2. For consistency, we proceed with eGP (M1) in the remainder of the paper.
Model M4, which omits XGBoost-derived covariates, relies solely on spatio-temporal effects in the latent Gaussian model. This leads to predictive distributions of council $s$ that are invariant at time point $t,\;t+12,\;t+24,\;\ldots$ in the test set (though the test set we used does not cover the whole cycle due to its 12-month length), reflecting purely seasonal and spatial patterns. As a result, its performance declines markedly, especially in metrics that weigh all events equally, such as AUC, CRPS, and unweighted $r^{C}$ and $r^{B}$. This highlights the importance of incorporating dynamic, forecast-driven covariates for accurate wildfire forecast.

### v2

@@ 3.2.2 Output Generation
We adopt a time-adjusted Super Learner cross-validation (CV) scheme (Van der Laan et al., 2007) to tune hyperparameters and generate one-month-ahead forecasts of fire counts and burnt area. Standard $k$-fold CV is unsuitable for time series, as random splits may allow training on observations that occur after the validation set, thereby introducing look-ahead bias. To avoid this, we employ a chronological CV strategy, in which the model is trained exclusively on past folds and evaluated on subsequent future folds.
Formally, let $D_{k},k=1,\ldots,12$ denote the subsets of records corresponding to calendar years 2011–2022 in the training set, and let $k_{0}$ mark the end of the warm-up period required to accumulate sufficient history for stable forecasting in the first validation fold. For each validation year $k_{\mathrm{val}}>k_{0}$, we fit the model on $\bigcup_{k<k_{\mathrm{val}}}D_{k}$ and evaluate it on $D_{k_{\mathrm{val}}}$. Hyperparameters are selected by optimising metrics computed on pooled out-of-fold predictions, i.e. aggregating predictions from all validation years $k_{\mathrm{val}}>k_{0}$} and evaluating the metric once on this combined set. After selection, we (i) refit on $\bigcup_{k<k_{\mathrm{val}}}D_{k}$ to generate within-training one-month-ahead forecasts for $D_{k_{\mathrm{val}}}$, and (ii) refit on $\bigcup_{k=1}^{12}D_{k}$ to forecast the test period (2023). This procedure prevents information leakage and provides performance estimates that reflect out-of-time generalisation, as illustrated in Figure 5.
Figure 5: Time-series–aware cross-validation is implemented using yearly blocks for both hyperparameter tuning and the generation of one-month-ahead forecasts. The first three years are excluded from the training set because feature construction based on a 36-month rolling window results in missing values during this period.
@@ 3.3 Stage II: Bayesian Latent Gaussian Modelling
We incorporate the one-month-ahead forecasts of fire count and burnt area from equations (5.3) and (2) as external covariates in a Bayesian latent Gaussian model, which yields the final probabilistic predictions. This two-stage setup reflects the principles of stacked generalisation (Wolpert, 1992), where the predictions from one model serve as informative inputs to a second, more interpretable model to enhance overall performance.
@@ 3.3.1 Model Structure and INLA
and this joint posterior could be approximated using MCMC. However, in many spatial applications, the marginal posterior distributions $\pi(u_{i}\mid\boldsymbol{y})$ and $\pi(\theta_{j}\mid\boldsymbol{y})$ are of greater practical interest than the full joint posterior, and knowledge of these marginals is sufficient for subsequent inference and analysis. Rue et al. (2009) proposed a fast approximation algorithm for marginal posteriors in latent Gaussian models using the integrated nested Laplace approximation (INLA). This algorithm was later reformulated to accommodate big data settings, enhancing numerical scalability and enabling fast inference (Van Niekerk et al., 2023). These developments form the foundation of the R-INLA package (Martins et al., 2013).
@@
with integration points $\boldsymbol{\theta}_{k}$ and weights $\Delta_{k}$. The marginal posterior of the linear predictors $\pi(\eta_{i}\mid\boldsymbol{y})$ is derived in a similar manner. Staring with the Gaussian approximation $\widetilde{\pi}(\boldsymbol{u}\mid\boldsymbol{\theta},\boldsymbol{y})$, the conditional posterior $\pi(\eta_{i}|\boldsymbol{\theta},\boldsymbol{y})$ is then also approximated as Gaussian. Its mean and variance can be efficiently computed leveraging the linear relationship $\boldsymbol{\eta}=\boldsymbol{Au}$ and some tricks on computing the $i$-th diagonal element of the inverse of the precision matrix associated with $\widetilde{\pi}(\boldsymbol{u}|\boldsymbol{\theta},\boldsymbol{y})$ (Van Niekerk et al., 2023). The marginal posterior $\pi(\eta_{i}|\boldsymbol{y})$ is then approximated by:
@@
Finally, Van Niekerk et al. (2023); van Niekerk and Rue (2024) proposed a low-rank correction to the mean of $\widetilde{\pi}(\boldsymbol{u}\mid\boldsymbol{\theta},\boldsymbol{y})$ using variational Bayes, further improving the approximation of both $\widetilde{\pi}(u_{i}\mid\boldsymbol{y})$ and $\widetilde{\pi}(\eta_{i}\mid\boldsymbol{y})$.
@@ 3.3.2 Likelihood Specification
Although Poisson and Tweedie likelihoods perform well in modelling general wildfire trends via XGBoost, they are not optimal in the latent Gaussian framework. Specifically, the Tweedie distribution, being composed of i.i.d. Gamma components, features a light right tail and is therefore unsuitable for capturing extreme burnt areas. Instead, we adopt the extended generalised Pareto (eGP) distribution (Naveau et al., 2016) for modelling the burnt area. This distribution combines features of Gamma and Pareto distributions by applying a power transformation to the standard Pareto distribution. It is defined on the positive real line and behaves like a Gamma for small values and like a Pareto for large values. The right tail behaviour is controlled by a shape parameter $\xi$, allowing flexibility in modelling heavy tails.
@@
We define three linear predictors for each council $s$ at time $t+1$: $\eta^{Z}_{s,t+1}$, $\eta^{C}_{s,t+1}$ and $\eta^{B}_{s,t+1}$, corresponding to fire presence ($Z$), fire count ($C$) and burnt area ($B$), respectively. The full hierarchical model is described as:
$Z_{s,t+1}\mid\eta_{s,t+1}^{Z}$ $\sim\text{Bernoulli}\{\text{logit}^{-1}(\eta_{s,t+1}^{Z})\}$
$\{Y_{s,t+1}^{C}\mid\eta_{s,t+1}^{C},Z_{s,t+1}=1\}$ $\sim\text{Trucated Poisson}\{\exp(\eta_{s,t+1}^{C})\}$
$\left\{\sqrt{Y_{s,t+1}^{B}}\mid\eta_{s,t+1}^{B},Z_{s,t+1}=1\right\}$ $\sim\text{eGP}\{\sigma\left(\eta_{s,t+1}^{B}\right),\xi,\kappa\}$
@@ 3.3.3 Effects in the Linear Predictor
Although the XGBoost predictions $\widehat{y}^{C}_{s,t+1}$ and $\widehat{y}^{B}_{s,t+1}$ could be included as fixed effects, doing so would impose a linear relationship with the linear predictor. Given the complex spatial and spatio-temporal dependencies present in the data, such a restriction is overly limiting. Instead, we treat the XGBoost outputs as smooth random effects, allowing the second-stage model to flexibly recalibrate the machine-learning predictions in a spatially coherent and uncertainty-aware manner. Specifically, we model the effect of $\widehat{y}^{(\cdot)}_{s,t+1}$ using a first-order random walk (RW1) prior on 20 discretised bins of the prediction $\widehat{y}_{s,t+1}^{(\cdot)}$:
$R_{k}-R_{k-1}\sim\mathcal{N}(0,\tau^{-1}_{R}),$ (6)
where $R_{k}$ represents the effect of the $k$-th bin of the discretised covariate and $\tau_{R}$ is the precision parameter. As the fire count and burnt area contribute to the Poisson and eGP likelihoods only when fire is present (i.e., when $y^{C}_{s,t+1}>0$) in the hurdle model training, we discretise $\widehat{y}_{s,t+1}^{C}$ and $\widehat{y}_{s,t+1}^{B}$ conditional on ${y}_{s,t+1}^{C}>0$ when constructing the $R_{k}$ in their respective linear predictors $\eta_{s,t+1}^{C}$ and $\eta_{s,t+1}^{B}$. By contrast, $\widehat{y}_{s,t+1}^{C}$ and $\widehat{y}_{s,t+1}^{B}$ are discretised unconditionally when constructing the corresponding $R_{k}$ in $\eta_{s,t+1}^{Z}$.
We incorporate spatially structured and unstructured effects through the Besag–York–Mollié model, using the reparameterised BYM2 formulation (Riebler et al., 2016; Simpson et al., 2017), as implemented in R-INLA. A BYM2 effect $\boldsymbol{b}$ combines a scaled intrinsic conditional autoregressive (CAR) component $\boldsymbol{\delta}$ (with unit variance) and unstructured noise $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$. Here, $\boldsymbol{I}$ represents the identity matrix with dimension defined by the length of $\boldsymbol{\epsilon}$. The effect is defined as
@@
[…]
$\eta_{s,t+1}^{C}$ $=\beta_{0}^{C}+\beta_{1}^{C}$ $G_{c}(s,t+1;\tau_{G_{c}},\phi_{G_{c}})+\beta_{2}^{C}$ $G_{d}(s,t+1;\tau_{G_{d}},\phi_{G_{d}})+T(t+1;\tau_{T}^{C})+R(\widehat{y}_{s,t+1}^{C};\tau_{R}^{C}),$
$\eta_{s,t+1}^{B}$ $=\beta_{0}^{B}+\beta_{1}^{B}$ $G_{c}(s,t+1;\tau_{G_{c}},\phi_{G_{c}})+\beta_{2}^{B}$ $G_{d}(s,t+1;\tau_{G_{d}},\phi_{G_{d}})+T(t+1;\tau_{T}^{B})+R(\widehat{y}_{s,t+1}^{B};\tau_{R}^{B}).$
@@ 3.3.4 Priors
We now elaborate on the priors used for the eGP likelihood hyperparameters ($\kappa$ and $\xi$) as well as those associated with the linear predictor. For $\kappa$ and $\xi$, we adopt Penalised Complexity (PC) priors following the framework of Simpson et al. (2017), which provide a principled mechanism to control model complexity by penalising deviations from a simpler base model. This deviation is measured via the Kullback–Leibler divergence (KLD; Kullback and Leibler 1951), and the corresponding distance is defined as $d=\sqrt{2\text{KLD}}$. An exponential prior is then placed on this distance, yielding a memoryless penalty on increasing model complexity. The PC prior for each parameter is obtained by transforming this exponential prior back to the original parameter scale.
In the special case of the standard GDP ($\kappa=1$), Opitz et al. (2018) derived a PC prior for $\xi>0$, using the base model with $\xi=0$. They computed the KLD between a GPD density $f_{\xi}(y)=f_{\text{GPD}}(y;\xi)$ and the base model $f_{\xi_{0}}(y)=f_{\text{GPD}}(y;\xi=0)$ as:
@@
where $\lambda$ is the rate parameter of the exponential distribution, controlling the strength of penalisation. The two priors are nearly indistinguishable for large $\lambda$ (e.g., $\lambda>3$), but diverge notably for smaller values, as illustrated in Figure 3 of Opitz et al. (2018).
@@
with derivation provided in the Supplementary Materials A.1. Following the approach for deriving a PC prior for the parameter $\xi$, we may either use the exact KLD in (10), or apply a second-order Taylor expansion around $\kappa=1$, which gives
@@
Figure 6 displays $\pi_{1}(\kappa)$ and $\pi_{2}(\kappa)$ under various values of $\lambda$. When $\lambda$ is large (e.g. $\lambda>5$), both priors concentrates around $\kappa=1$, exhibiting similar behaviour. As $\lambda$ decreases, the mode of $\pi_{1}(\kappa)$ shifts leftwards towards zero and diverge as $\kappa\rightarrow 0$, whereas $\pi_{2}(\kappa)$ remains locally symmetric around $\kappa=1$. This suggests that $\pi_{1}(\kappa)$ always shrinks $\kappa$ towards a value in $(0,1]$, and it is less suitable for expressing weakly informative priors over $\kappa>1$ compared to $\pi_{2}(\kappa)$. To determine the more appropriate prior between $\pi_{1}(\kappa)$ and $\pi_{2}(\kappa)$, we consider the interpretation and functional role of $\kappa$. According to Naveau et al. (2016), $\kappa$ governs the lower-tail behaviour of the cumulative distribution function of an eGP random variable $Y$ via
@@
Figure 6: PC priors for $\kappa$ based on exact KLD in (10) (left) and the approximated KLD around $\kappa=1$ (11) (right) under penalisation rates $\lambda\in\{0.1,0.5,1,2,5,10\}$.
@@
Here, we adopt Gamma priors rather than PC priors for the parameters in $T(\cdot)$ and $R(\cdot)$ because numerical instabilities were observed during model fitting across a wide range of PC prior specifications, from weakly informative to more informative choices. These instabilities were substantially mitigated after switching to Gamma priors.
@@ 4.1 Model Comparison
To evaluate the effectiveness of incorporating XGBoost-based wildfire forecasts into our modelling framework and to assess the appropriateness of the eGP likelihood for wildfire modelling, we compare several variations of the latent Gaussian model differing in their likelihood choices and linear predictor components. Let M1 denote the full two-stage model introduced in Section 3. To examine the role of the eGP likelihood, we substitute it with two alternative distributions commonly used in environmental applications: the Gamma and Weibull likelihoods, yielding models M2 and M3, respectively. Additionally, we construct M4 to emulate a setting in which future environmental covariates are unavailable. Specifically, the XGBoost-derived random effects are replaced by lag-1 values of the FWI and temperature. We retain FWI because it is one of the most widely used fire danger indicators, and temperature due to its high feature importance, as shown in Figure 8. All remaining random effects are kept identical to those in M1.
We use wildfire data from 2011 to 2022 as the training set and data from 2023 as the test set. One-month-ahead forecasts of fire count and burnt area are generated as described in Section 3.2.2. The posterior predictive distributions of the fire presence $Z_{s,t+1}$, fire count $Y_{s,t+1}^{C}$ and square-root-transformed burnt area $\sqrt{Y_{s,t+1}^{B}}$ are obtained from 1000 posterior simulations. To evaluate performance, we use Area Under Curve (AUC) to assess the predictive accuracy for fire presence, where the posterior mean of $\widehat{Z}_{s,t+1}$ serves as the estimate of $\mathbb{P}(Z_{s,t+1}=1)$. The positive burnt area $\sqrt{\widehat{Y}_{s,t+1}^{B}}\mid Z_{s,t+1}=1$ is evaluated by continuous ranked probability score (CRPS), which is a proper scoring rule that measures the difference between a predictive distribution $F$ and a single observation $y$ by
$\text{CRPS}(F,y)=\int_{\mathbb{R}}\left[F(v)-\mathbbm{1}(v\geq y)\right]^{2}\mathrm{d}v,$
@@
Since forecasts are typically communicated as probabilities over categorised bins, with particular emphasis on large events, we also compute weighted binned scores following the structure proposed by Opitz (2023). Specifically, for a fire count threshold vector $\boldsymbol{h}^{C}=(0,1,2,\ldots,10,15,20,25,30)$, the weighted scoring for fire count $r^{C}$ is the weighted sum of squared residuals between predicted and empirical probabilities across all observations and thresholds:
$r^{C}=\sum_{s,t+1}\sum_{h\in\boldsymbol{h}^{C}}w^{C}(h)\left[\widehat{\mathbb{P}}(Y_{s,t+1}^{C}\leq h)-\mathbbm{1}(Y_{s,t+1}^{C}\leq h)\right]^{2},$
@@
which increases approximately linearly with $h$. Here, $\mathbb{P}(Y_{s,t+1}^{C}\leq h)$ is the unconditional probability of $Y_{s,t+1}^{C}$ obtained by integrating out $Z_{s,t+1}$ from the conditional distribution $Y_{s,t+1}^{C}\mid Z_{s,t+1}=1$.
@@
$r^{B}=\sum_{s,t+1}\sum_{h\in\boldsymbol{h}^{B}}w^{B}(h)\left[\widehat{\mathbb{P}}(Y_{s,t+1}^{B}\leq h)-\mathbbm{1}(Y_{s,t+1}^{B}\leq h)\right]^{2},$
@@
Metric M1 (eGP) M2 (Gamma) M3 (Weibull) M4 (eGP, lagged covariates)
AUC 0.862 0.857 0.860 0.815
CRPS 4.75 4.83 4.78 4.89
Unweighted $r^{C}$ 362 371 366 382
Weighted $r^{C}$ 4.84 4.93 4.88 5.14
Unweighted $r^{B}$ 565 578 577 594
Weighted $r^{B}$ 14.2 14.5 14.6 14.7
Table 1 summarises the model performance on the test set in 2023 across six evaluation metrics. Among models M1, M2, and M3, which differ only in the likelihood, the eGP likelihood achieves slightly better performance across all metrics. M1 and M3 perform comparably, as both eGP and Weibull likelihoods can accommodate heavy-tailed data through their shape parameter. In contrast, M2 employs a light-tailed distribution (Gamma) for burnt area and performs worst on five of the six metrics. Notice that although the likelihood is modified only for burnt area, changes also propagate to metrics such as AUC and binned scores for counts, owing to the shared spatio-temporal effects in the joint modelling of fire presence, fire count, and burnt area. Nonetheless, performance differences between M1, M2, and M3 are marginal, and the reasons for this will be further discussed in Section 5.2. For consistency, we proceed with eGP (M1) in the remainder of the paper.
Model M4, which excludes XGBoost-derived covariates, relies only on lagged temperature and FWI covariates together with spatio-temporal effects in the latent Gaussian model to produce one-month-ahead forecasts. A substantial deterioration in performance is observed across all six evaluation metrics, with particularly pronounced declines in metrics that weight all events equally, such as AUC, CRPS, and the unweighted correlations $r^{C}$ and $r^{B}$. These results underscore the importance of incorporating dynamic, forecast-driven covariates that capture complex spatio-temporal structure for accurate wildfire forecasting.